OpenStudy (anonymous):

what is the solution of each equation? 2 (h+8)-h=h+16

Parth (parthkohli):

Start by adding \(h\) to both sides. \( \color{Black}{\Rightarrow 2(h + 8 ) = 2h + 16 }\) \( \color{Black}{\Rightarrow 2h + 16 = 2h + 16 }\) As \(h\) applies to all real numbers, we may say: \( \color{Black}{\Rightarrow h \Large {\epsilon} \mathbb{R} }\)

OpenStudy (anonymous):

A.8 B.-8 C.infinitely many solutions D.No solution

Parth (parthkohli):

Infinitely many solutions.

Parth (parthkohli):

This is because you may substitute the value of \(h\) with any real number and the equation will satisy.

Parth (parthkohli):

satisfy*

OpenStudy (anonymous):

thank you